OpenStudy (hihi67):

6x/8y^3 ÷ (2x^2)^3/9x^4 =

OpenStudy (hihi67):

(im checking my answers)

OpenStudy (anonymous):

I got 54x^11/y^3 seems wrong but I'm not sure if I misinterpreted the inital question.

OpenStudy (hihi67):

i did this: 6x/8y^3 ÷ (2x^2)^3/9x^4 = 2x/2y^3 ÷ x^6/3y^4 = 2x/2y^3 * 3y/x^5 = 6xy/2x^5

OpenStudy (hihi67):

and also : (-x^2)^3/4z ÷ 3x/8z^2 =

OpenStudy (anonymous):